1. How long will a bus take to travel 150 km at an average speed of

40 km/h?

2.

2. A cheetah can run at 30 m/s, but only for about 12 s. How far

will it run in that time?

3. It takes a cheetah just 3 s to reach its top speed of 30 m/s.

What is its acceleration?



1
Expert's answer
2022-02-15T16:14:13-0500

Solution.

"1. l=150km;"

"v=40km\/h;"

"l=vt\implies t=\dfrac{l}{v}=\dfrac{150km}{40km/h}=3.75h."

"2. v=30 m\/s;"

"t=12s;"

"l=vt=30 m/s\sdot12s=360m;"

"3. v=30m\/s;"

"v_0=0m\/s;"

"t=3s;"

"a=\dfrac{v-v_0}{t}=\dfrac{30m/s}{3s}=10m/s^2;"

Answer: "1. l=3.75h;"

"2.l=360m;"

"3. a=10 m\/s^2."